# Thousand Voices of Trauma: A Large-Scale Synthetic Dataset for Modeling Prolonged Exposure Therapy Conversations

### v3

@@ Abstract
Thousand Voices of Trauma: A Large-Scale Synthetic Dataset for Modeling Prolonged Exposure Therapy Conversations
Suhas BN1  Dominik Mattioli1  Saeed Abdullah1 Rosa I. Arriaga2  Chris W. Wiese3  Andrew M. Sherrill4 1College of Information Sciences and Technology, Penn State University, USA 2School of Interactive Computing, Georgia Tech, USA 3School of Psychology, Georgia Tech, USA 4Department of Psychiatry and Behavioral Sciences, Emory University, USA {bnsuhas,saeed}@psu.edu, andrew.m.sherrill@emory.edu
@@ 1 Introduction
The intersection of mental health care and artificial intelligence presents unprecedented opportunities alongside significant challenges. AI system development faces particular obstacles in trauma-focused therapy due to the sensitive nature of patient experiences and strict privacy regulations, which make the collection of real-world data extremely challenging (Mathur et al., 2022).
Moreover, existing datasets frequently lack the diversity and clinical depth needed to train robust AI systems capable of serving diverse populations effectively Marr (2023). Prolonged Exposure (PE) therapy—an evidence-based treatment for post-traumatic stress disorder (PTSD) (Foa et al., 2007)—offers a structured therapeutic approach that could especially benefit from AI support.
However, current mental health conversation datasets are often too small (Zantvoort et al., 2024), lack demographic diversity (Sadeh-Sharvit, 2024), and do not capture the nuanced progression of trauma-focused therapy sessions (Mathur et al., 2022).
To address these gaps, we introduce Thousand Voices of Trauma, a synthetic benchmark dataset comprising 500 clinical sessions. Each session is divided into six distinct therapeutic phases aligned with core evaluation criteria of PE therapy. Each phase contains multiple therapist-client exchanges, which can be analyzed either independently or as part of the complete session flow, resulting in 3000 distinct clinical conversations. These phases, as defined by Foa et al. (2007), correspond to specific stages within the PE therapy session: a) Orientation to Imaginal Exposure, b) Imaginal Exposure Duration, c) Monitoring SUDS Ratings, d) Reinforcing Comments, e) Eliciting Thoughts and Feelings, and f) Processing the Imaginal.
@@
This dataset also addresses other real-world limitations that often hinder mental health research and safe AI model development. For example, while privacy concerns often restrict data access, synthetic data can circumvent typical ethical and legal barriers. It also overcomes other common issues with real-world data, such as incompleteness, inconsistency, and small sample sizes—especially among minority groups.
@@ 1.1 Key Contributions
    Scale and Diversity: Our dataset is one of the largest collections of synthetic trauma-focused therapy conversations, encompassing varied demographic profiles across age, gender, ethnicity, and cultural backgrounds. The use of synthetic data addresses ethical concerns related to patient privacy and consent, while its diversity supports the development of more inclusive and culturally competent AI systems.
@@
    Clinical Depth: Each conversation is grounded in evidence-based PE therapy principles and incorporates 20 distinct trauma types, 10 trauma-related behaviors, and 5 co-occurring conditions. This clinical specificity renders the dataset particularly valuable for specialized therapeutic applications, including clinician training.
@@
    Structured Evaluation Framework: The six-conversation structure per session enables a comprehensive examination of different interaction trajectories, spanning from the initial assessment to trauma processing and progress evaluation.
@@
    Baseline: We establish a baseline dataset for training AI systems in trauma-focused therapy, facilitating standardized comparisons across different approaches.
The remainder of this paper is structured as follows: Section 2 reviews related work in mental health. Section 3 describes the datasets and details the synthetic data generation methodology. Section 4 presents the expert evaluation. Section 5 introduces the benchmark setup and evaluation method. Sections 69 discuss future directions, data availability, ethical considerations, and limitations.
@@ 2 Related Work
Prolonged Exposure (PE) therapy, an evidence-based treatment for PTSD, relies on structured exposure to trauma-related narratives (Watkins et al., 2018). However, there is a lack of trained professionals who can provide PE therapy (Rauch et al., 2017, 2023). As a result, there is an urgent need for AI applications to support PE therapy delivery and training. This underscores the need for clinically valid and diverse datasets for AI development and evaluation. Large-scale language models (LLMs), like the GPT series, have shown potential in generating synthetic datasets that mimic human-like text, addressing challenges such as data scarcity and privacy concerns (Hämäläinen et al., 2023; Giuffrè and Shung, 2023; Zhu, 2024; Li et al., 2024). However, for applications like PE therapy—which require alignment with trauma-focused frameworks, diverse demographic representation, and strict ethical safeguards—current research still lacks tailored solutions.
@@
Synthetic datasets show promise in mental health applications, with studies exploring LLM-based data generation to address data scarcity. Wu et al. (Wu et al., 2023, 2024) introduced zero-shot and few-shot learning frameworks to augment PTSD diagnostic datasets, producing synthetic transcripts that outperform baselines. The latter work used role-prompting and structured prompts to create realistic synthetic clinical interviews.
@@
However, these approaches have not been specifically adapted to the structured narratives essential for PE therapy. Efforts to enhance demographic diversity in synthetic datasets are growing. Mori et al. (2024) and Lozoya et al. (2023) examined how synthetic data reflects demographic variation, highlighting biases in LLM outputs, especially regarding race and gender, and stressing fairness in mental health datasets. Techniques like patient vignette simulation (Denecke, 2024) and adaptive prompts for non-English contexts (e.g., SAPE for Spanish (Lozoya et al., 2024)) show early progress toward inclusivity. However, trauma-type diversity and PE-specific scenarios remain unexplored. Additionally, Chen et al. (2024) underscored the need for systematic benchmarking using metrics like F1-score, AUC, and balanced accuracy, but these tools have not yet been applied to datasets focused on diverse trauma types or PE therapy.
@@
Privacy and ethics are central to synthetic dataset generation, with studies focusing on privacy-preserving methods. Recent works (Chuang et al., 2024a; Meoni et al., 2024; Suhas et al., 2023; BN and Abdullah, 2022; Chuang et al., 2024b) highlight privacy-preserving machine learning, protected health information (PHI) exclusion, and semantic filtering to maintain privacy compliance while preserving data utility. While these methods offer strong safeguards for general clinical use, ethical risks—such as generating harmful trauma narratives or victim-blaming—remain underexplored, especially in sensitive contexts like trauma-focused therapies such as PE.
@@
While synthetic dataset generation using LLMs (Wu et al., 2023, 2024; Xu et al., 2024) has advanced in addressing general clinical challenges, gaps in the literature remain for PE therapy. These include the lack of trauma-type diversity, limited demographic inclusivity evaluation, insufficient alignment with frameworks like DSM-5 PTSD criteria, and underdeveloped ethical safeguards specific to trauma-focused contexts. Synthetic datasets can also help mitigate representation biases in AI models for mental healthcare delivery.
For example, American Psychiatric Association (2013) reported that non-Hispanic White adults (25.0%) were more likely to receive mental health services than non-Hispanic Black (18.3%), Hispanic (17.3%), and Asian (13.9%) adults. Including such underrepresented groups in synthetic datasets might partially address the training data gap. This paper seeks to bridge these gaps by exploring the existing knowledge base and identifying pathways to tailor synthetic data generation for PE therapy applications.
Figure 1: The figure presents the demographic distribution of generated participant profiles across four categories: gender, age, ethnicity, and relationship status. The gender distribution shows a nearly balanced representation of males (247) and females (222), with a smaller group identifying as non-binary (31) mimicking real world gender distributions U.S. Census Bureau (2025). The Age Distribution spans from under 10 to over 90 years, with most participants between 30 and 70 years. The Ethnicity Distribution highlights diverse backgrounds, with the largest groups being Latin American (70), North American (67), South East Asian (64), South Asian (63), and European (63), followed by Oceania (62), Middle Eastern (59), and African (52) participants. The Relationship Status chart reveals that most participants are married (180) or single (149), while smaller numbers are in relationships (70), divorced (51), or widowed (50).
Figure 2: The figure presents the distribution of exhibited behaviors and types of trauma among participants. The Exhibited Behaviors chart shows that nightmares (117 participants) are the most common symptom, followed by substance abuse (104), compulsive behaviors (101), aggression (100), and hypervigilance (99). Other frequently reported behaviors include self-blame (97), withdrawal (95), flashbacks (95), avoidance (91), and dissociation (83). The Types of Trauma chart highlights that witnessing violence (53 participants) and bullying (51) are the most frequently reported traumatic experiences. Other significant trauma types include neglect (45), medical trauma (43), childhood trauma (42), and the loss of a loved one (39). Natural disasters, sexual abuse, physical abuse, and emotional abuse were each reported by 39 to 38 participants, while combat or war experiences and accidents were each cited by 36 participants.
@@ 3.1 Dataset: Simulated Therapy Session Profiles
To systematically examine trauma narratives and therapeutic dynamics, we constructed a synthetic dataset comprising 3,000 simulated therapy conversations. These dialogues were generated using Claude Sonnet 3.5 (Anthropic, 2024), selected as the foundational model for the Thousand Voices of Trauma dataset based on its demonstrated ability to produce coherent, contextually grounded, and richly detailed dialogue during preliminary assessments. The rationale for prioritizing Claude Sonnet 3.5 was centered on its superior generative capacity, which allowed the efficient production of high-quality conversations that aligned closely with the principles of prolonged exposure therapy (PE). This choice allowed us to pay more attention to the methodological challenges of scenario construction, prompt engineering, and the integration of PTSD-specific symptomatology and therapeutic constructs.
@@
Although comparative evaluation of other large language models remains a compelling direction for future investigation, the initial deployment of a robust, high-performing model was critical to establishing a reliable and practically useful foundation for trauma-focused research. This decision facilitated the immediate applicability of the dataset for downstream tasks, including the modeling of trauma-informed therapeutic interactions and the development of evaluation frameworks for clinical AI tools.
@@
Each simulated profile integrates structured metadata, including client demographics (e.g., age, gender, living situation), therapist attributes, and session-level variables such as trauma type, therapeutic phase, and discussion topics. To ensure diversity and ecological validity, both deterministic (rule-based) and probabilistic (sampling-based) generation methods were used in constructing the scenarios.
To illustrate model-specific stylistic and structural variations, we provide a comparative set of examples generated by multiple state-of-the-art Frontier Models. Full prompt templates and representative conversations for each phase of the therapeutic protocol are included in Appendix A and B.
3.2 Client Profile Generation
@@
Client profiles included age, gender, relationship status, occupation, living situation, and ethnicity. Ages ranged from 18 to 80, divided into six groups: 18-30, 31-40, 41-50, 51-60, 61-70, 71-80. We assigned gender using weighted probabilities: 50% male, 49% female, and 1% non-binary U.S. Census Bureau (2025). Relationship status, occupation, and living situation were age-specific—for example, clients aged 20-30 were more likely to be “Single,” “Student,” and “With parents,” while those 60-70 were often “Widowed,” “Retired,” and “Alone.” A validation function ensured logical consistency. We randomly assigned ethnicity from eight global regions: South Asian, Middle Eastern, African, North American, Oceanian, European, South East Asian, and Latin American. We assigned co-occurring conditions with weighted probabilities Bilevicius et al. (2020); Jennifer et al. (2024); Hagiwara et al. (2022): None (25%), Anxiety (25%), Depression (30%), Substance Use Disorder (10%), and Chronic Pain (10%). We also assigned clients 1–3 trauma-related behaviors from ten options, including avoidance, hypervigilance, flashbacks, nightmares, self-blame, substance abuse, aggression, withdrawal, dissociation, and compulsive behaviors. The options represent a range of cognitive, emotional, and behavioral responses typically associated with trauma, which aligns with trauma-informed care principles (SAMHSA, 2015, 2014).
3.3 Therapist Profile Generation
@@
We generated therapist profiles with ages ranging from 25 to 65, divided into four age groups: 25-34, 35-44, 45-54, and 55-65. We assigned therapist gender using the same weighted probabilities as client gender.
3.4 Therapy Context Generation
@@
To create varied therapeutic contexts, we defined both a broad trauma type representing the general category of traumatic experience, and a more specific session topic providing the focus for the simulated interaction. We generated the therapy context by randomly selecting a trauma type from twelve possibilities (SAMHSA, 2015, 2014): physical abuse, emotional abuse, sexual abuse, neglect, natural disasters, accidents, combat or war experiences, loss of a loved one, witnessing violence, bullying, childhood trauma, and medical trauma. A session topic was also randomly selected from twenty possibilities, such as car accident trauma, workplace trauma, domestic violence, natural disaster experience, military combat experience, loss of a loved one, severe illness, divorce or relationship breakup, witnessing violence, racial trauma, and refugee experiences. This independent random selection of trauma type and session topic allows for a wide range of generated scenarios. While this methodology can result in pairings without an immediate thematic link (e.g., ‘natural disaster’ type with ‘workplace trauma’ topic), such combinations can also reflect the complexity observed in real-world therapy, where session discussions may focus on various life stressors influenced by, or co-existing with, the client’s primary trauma history.
3.5 Session Profile Assembly
@@
Each complete session profile combined a validated client profile, a therapist profile, and the generated therapy context, including the trauma type and session topic.
3.6 Dataset Statistics
The dataset comprises 500 simulated participants (ages 18-80 years, M = 49.3). The gender distribution includes 247 male (49.4%), 222 female (44.4%), and 31 non-binary (6.2%) participants.
@@ 3.6 Dataset Statistics
The ethnicity distribution is as follows: Latin American (70, 14.0%), North American (67, 13.4%), South East Asian (64, 12.8%), South Asian (63, 12.6%), European (63, 12.6%), Oceania (62, 12.4%), Middle Eastern (59, 11.8%), and African (52, 10.4%). Regarding relationship status, participants were predominantly married (180, 36.0%) or single (149, 29.8%), with others reporting being in a relationship (70, 14.0%), divorced (51, 10.2%), or widowed (50, 10.0%). See Figure 1 & 2 for more details.
@@
Figure 5: Stacked Bar charts illustrating therapist ratings (N=7) across four key dimensions of the synthetic PE therapy sessions—Content Depth, Perceived Value, Session Appropriateness, and Patient Engagement.
Figure 6: Violin plots illustrating the distribution of emotional intensity scores across six therapy stages scored by Claude Sonnet 3.5. Emotional intensity generally starts at moderate levels during orientation, peaks during exposure and monitoring, and then tapers off through reinforcing, reflection, and processing phases.

### v4

@@ 1 Introduction
The intersection of mental health care and artificial intelligence presents unprecedented opportunities alongside significant challenges. AI system development faces particular obstacles in trauma-focused therapy due to the sensitive nature of patient experiences and strict privacy regulations, which make the collection of real-world data extremely challenging [1].
Moreover, existing datasets frequently lack the diversity and clinical depth needed to train robust AI systems capable of serving diverse populations effectively [2]. Prolonged Exposure (PE) therapy—an evidence-based treatment for post-traumatic stress disorder (PTSD) [3]—offers a structured therapeutic approach that could especially benefit from AI support.
However, current mental health conversation datasets are often too small [4], lack demographic diversity [5], and do not capture the nuanced progression of trauma-focused therapy sessions [1].
We introduce Thousand Voices of Trauma, a synthetic benchmark dataset comprising 500 clinical sessions, each structured into six core phases of Prolonged Exposure (PE) therapy. These phases, based on Foa et al. [3], span the full therapeutic arc: (a) Orientation to Imaginal Exposure, (b) Imaginal Exposure Duration, (c) Monitoring SUDS Ratings, (d) Reinforcing Comments, (e) Eliciting Thoughts and Feelings, and (f) Processing the Imaginal. Each phase includes multiple therapist–client exchanges, which can be analyzed independently or as a complete session flow, yielding 3,000 structured clinical conversations. All dialogues were generated using Sonnet 3.5 [6], guided by clinically informed prompts to ensure alignment with PE structure and therapeutic fidelity.
@@
This dataset also addresses real-world limitations often hindering mental health research and safe AI model development. For example, while privacy concerns usually restrict data access, synthetic data can circumvent typical ethical and legal barriers. It also overcomes other common issues with real-world data, such as incompleteness, inconsistency, and small sample sizes, especially among minority groups.
@@ 1.1 Key Contributions
    Scale and Diversity: To our knowledge, this is the first large-scale structured dataset of therapy conversations grounded in PTSD treatment protocols, covering diverse demographics across age, gender, ethnicity, and culture. Synthetic generation mitigates privacy concerns and promotes inclusive, culturally aware AI development.
@@
    Clinical Depth: Grounded in evidence-based PE therapy, the dataset spans 20 trauma types, 10 trauma-related behaviors, and 5 co-occurring conditions—supporting use in clinician training and specialized therapeutic applications.
@@
    Structured Evaluation Framework: Each session includes six conversations, enabling analysis of interaction trajectories from intake to trauma processing and progress evaluation.
@@
    Baseline Resource: Provides a standardized reference for training and evaluating AI models in trauma-focused therapy.
The rest of the paper is organized as follows: Section 2 reviews related work; Section 3 details data generation; Section 4 presents expert evaluation; Section 5 outlines the benchmark; and Sections 69 discuss future work, data availability, ethics, and limitations.
@@ 2 Related Work
Prolonged Exposure (PE) therapy, an evidence-based treatment for PTSD, relies on structured exposure to trauma-related narratives [7]. However, there is a lack of trained professionals who can provide PE therapy [8, 9]. As a result, there is an urgent need for AI applications to support PE therapy delivery and training. This underscores the need for clinically valid and diverse datasets for AI development and evaluation. Large-scale language models (LLMs), like the GPT series, have shown potential in generating synthetic datasets that mimic human-like text, addressing challenges such as data scarcity and privacy concerns [10, 11, 12, 13]. However, for applications like PE therapy—which require alignment with trauma-focused frameworks, diverse demographic representation, and strict ethical safeguards—current research still lacks tailored solutions.
@@
Synthetic datasets show promise in mental health applications, with studies exploring LLM-based data generation to address data scarcity. Wu et al. [14, 15] introduced zero-shot and few-shot learning frameworks to augment PTSD diagnostic datasets, producing synthetic transcripts that outperform baselines. The latter work used role-prompting and structured prompts to create realistic synthetic clinical interviews.
@@
However, these approaches have not been specifically adapted to the structured narratives essential for PE therapy. Efforts to enhance demographic diversity in synthetic datasets are growing. Mori et al. [16] and Lozoya et al. [17] examined how synthetic data reflects demographic variation, highlighting biases in LLM outputs, especially regarding race and gender, and stressing fairness in mental health datasets. Techniques like patient vignette simulation [18] and adaptive prompts for non-English contexts (e.g., SAPE for Spanish [19]) show early progress toward inclusivity. However, trauma-type diversity and PE-specific scenarios remain unexplored. Additionally, Chen et al. [20] underscored the need for systematic benchmarking using metrics like F1-score, AUC, and balanced accuracy, but these tools have not yet been applied to datasets focused on diverse trauma types or PE therapy.
@@
Privacy and ethics are central to synthetic dataset generation, with studies focusing on privacy-preserving methods. Recent works [21, 22, 23, 24, 25] highlight privacy-preserving machine learning, protected health information (PHI) exclusion, and semantic filtering to maintain privacy compliance while preserving data utility. While these methods offer strong safeguards for general clinical use, ethical risks—such as generating harmful trauma narratives or victim-blaming—remain underexplored, especially in sensitive contexts like trauma-focused therapies such as PE.
@@
While synthetic dataset generation using LLMs [14, 15, 26] has advanced in addressing general clinical challenges, gaps in the literature remain for PE therapy. These include the lack of trauma-type diversity, limited demographic inclusivity evaluation, insufficient alignment with frameworks like DSM-5 PTSD criteria, and underdeveloped ethical safeguards specific to trauma-focused contexts. Synthetic datasets can also help mitigate representation biases in AI models for mental healthcare delivery. For example, American Psychiatric Association [27] reported that non-Hispanic White adults (25.0%) were more likely to receive mental health services than non-Hispanic Black (18.3%), Hispanic (17.3%), and Asian (13.9%) adults. Including such underrepresented groups in synthetic datasets might partially address the training data gap. This paper seeks to bridge these gaps by exploring the existing knowledge base and identifying pathways to tailor synthetic data generation for PE therapy applications.
Figure 1: Demographic distribution of synthetic participants across gender, age, ethnicity, and relationship status. Most identified as male (247) or female (222), with 31 non-binary participants [28]. Ages spanned under 10 to over 90, with a majority between 30–70. Ethnicities were diverse, led by Latin American, North American, and South/Southeast Asian groups. Most participants were married or single.
Figure 2: Distribution of trauma types and exhibited behaviors in synthetic participants. Nightmares, substance abuse, and compulsive behaviors were most common. Top trauma types included witnessing violence, bullying, neglect, and medical trauma. Less frequent but notable were abuse-related and combat-related experiences.
@@ 3.1 Dataset: Simulated Therapy Session Profiles
To systematically examine trauma narratives and therapeutic dynamics, we constructed a synthetic dataset comprising 3,000 simulated therapy conversations. These dialogues were generated using Claude Sonnet 3.5 [6], selected as the foundational model for the Thousand Voices of Trauma dataset based on its demonstrated ability to produce coherent, contextually grounded, and richly detailed dialogue during preliminary assessments. The rationale for prioritizing Claude Sonnet 3.5 was centered on its superior generative capacity, which allowed the efficient production of high-quality conversations that aligned closely with the principles of prolonged exposure therapy (PE). This allowed us to focus on methodological challenges such as scenario construction, prompt design, and fidelity modeling.
@@
Although comparing other large-language models remains a valuable future direction, deploying a strong initial model was essential for establishing a reliable foundation for trauma-focused research. This enabled immediate use of the dataset for downstream tasks like modeling therapeutic interactions and developing evaluation frameworks for clinical AI.
@@
Each simulated profile integrates structured metadata, including client demographics (e.g., age, gender, living situation), therapist attributes, and session-level variables such as trauma type, therapeutic phase, and discussion topics. To ensure diversity and ecological validity, both deterministic (rule-based) and probabilistic (sampling-based) generation methods were used in constructing the scenarios. To illustrate model-specific stylistic and structural variations, we provide a comparative set of examples generated by multiple state-of-the-art Frontier Models. Full prompt templates and representative conversations are included in Appendix A and B.
3.2 Session Design and Composition
@@
Client Profile Generation: Client profiles included age, gender, relationship status, occupation, living situation, and ethnicity. Ages ranged from 18 to 80, divided into six groups: 18-30, 31-40, 41-50, 51-60, 61-70, 71-80. We assigned gender using weighted probabilities: 50% male, 49% female, and 1% non-binary U.S. Census Bureau [28]. Relationship status, occupation, and living situation were age-specific—for example, clients aged 20-30 were more likely to be “Single,” “Student,” and “With parents,” while those 60-70 were often “Widowed,” “Retired,” and “Alone.” A validation function ensured logical consistency. We randomly assigned ethnicity from eight global regions: South Asian, Middle Eastern, African, North American, Oceanian, European, South East Asian, and Latin American. We assigned co-occurring conditions with weighted probabilities Bilevicius et al. [29], Jennifer et al. [30], Hagiwara et al. [31]: None (25%), Anxiety (25%), Depression (30%), Substance Use Disorder (10%), and Chronic Pain (10%). We also assigned clients 1–3 trauma-related behaviors from ten options, including avoidance, hypervigilance, flashbacks, nightmares, self-blame, substance abuse, aggression, withdrawal, dissociation, and compulsive behaviors. The options represent a range of cognitive, emotional, and behavioral responses typically associated with trauma, which aligns with trauma-informed care principles [32, 33].
@@
Therapist Profile Generation: We generated therapist profiles with ages ranging from 25 to 65, divided into four age groups: 25-34, 35-44, 45-54, and 55-65. We assigned therapist gender using the same weighted probabilities as client gender.
@@
Therapy Context Generation: To generate diverse therapeutic scenarios, each session paired a broad trauma type with a more specific session topic. Trauma types were randomly selected from twelve categories [32, 33], including physical, emotional, or sexual abuse; neglect; natural disasters; accidents; combat; bereavement; witnessing violence; bullying; childhood trauma; and medical trauma. Session topics were independently chosen from twenty possibilities, such as car accidents, domestic violence, workplace trauma, natural disasters, military combat, loss of a loved one, severe illness, divorce, racial trauma, and refugee experiences. This independent sampling supports a wide range of combinations. While some pairings may seem loosely linked (e.g., ‘natural disaster’ type with ‘workplace trauma’ topic), they reflect real-world therapy dynamics where discussions often explore co-occurring stressors or secondary experiences shaped by the client’s trauma history.
@@
Session Profile Assembly: Each complete session profile combined a validated client profile, a therapist profile, and the generated therapy context, including the trauma type and session topic.
@@
Dataset Statistics: The dataset comprises 500 simulated participants (ages 18-80 years, M = 49.3). The gender distribution includes 247 male (49.4%), 222 female (44.4%), and 31 non-binary (6.2%) participants. The ethnicity distribution is as follows: Latin American (70, 14.0%), North American (67, 13.4%), South East Asian (64, 12.8%), South Asian (63, 12.6%), European (63, 12.6%), Oceania (62, 12.4%), Middle Eastern (59, 11.8%), and African (52, 10.4%). Regarding relationship status, participants were predominantly married (180, 36.0%) or single (149, 29.8%), with others reporting being in a relationship (70, 14.0%), divorced (51, 10.2%), or widowed (50, 10.0%). See Figure 1 & 2 for more details.
@@
Figure 5: Therapist ratings ($N{=}7$) across four dimensions of synthetic PE sessions: Content Depth, Perceived Value, Session Appropriateness, and Patient Engagement.
Figure 6: Violin plots showing distribution of emotional intensity scores across six therapy stages as rated by Claude Sonnet 3.5. Intensity starts at moderate levels during orientation, peaks during exposure and monitoring, and then tapers off through reinforcing, reflection, and processing phases.